OpenStudy (anonymous):

Solve the following system of equations. 2x – 3y + z = –1 3x + 2y + 2z = –1 x – y – 3z = –4 (–1, 0, 1) (1, 1, 4) (–1, –1, –2) (1, 2, 1)

OpenStudy (anonymous):

Would i use elimination or substitution ?

OpenStudy (anonymous):

well even i have never solved such a qn before so let me have a go - 2x – 3y + z = –1 3x + 2y + 2z = –1 x – y – 3z = –4 Now in eqn 3, x-y-3z = -4 y = x-3z+4 using this value of y in eqn 1, 2x - 3(x-3z+4) + z = -1 2x - (3x-9z+12) + z = -1 2x-3x+9z-12+z+1 = 0 i.e. 10z-x-11 = 0 Now using same value of y in eqn 2, 3x+2y+2z = -1, 3x + 2(x-3z+4) + 2z = -1, 3x+2x-6z+8+2z+1 = 0 5x-4z+9 = 0 so x-10z =-11 and 5x-4z = 9 i.e. 5x-50z = -55 and 5x-4z =- 9 i.e. -46z =- 46 or z = 1. now 5x-4z = 9 5x-4 = -9 5x = -5 hence x = -1 Now using x = -1, z = 1 in y =x-3z+4, y = -1 -3.1 + 4 => -4+4 or y = 0 So solution set is (-1,0,1) i.e. A i have done this using substitution of y in terms of other 2 variables.

OpenStudy (anonymous):

oh wow you made it look easy lol thanks(:!